OpenStudy (kendricklamar2014):

List three different ways to write 5^11 as the product of two powers. Explain why all three of your expressions are equal to 5^11.

OpenStudy (kendricklamar2014):

@Shay17

OpenStudy (anonymous):

5^11 is like multiplying 5 by itself 11 times

OpenStudy (anonymous):

So it can be similar to \[5^3 \times 5^8 \] Since that also says that 5 is multiplied 11 times by itself

OpenStudy (anonymous):

5^7(5^4) 5^1(5^10) 5^6(5^5)

OpenStudy (anonymous):

Now that you have an example and explanation you can make another way and explain why that would be the same

OpenStudy (anonymous):

All three expressions are 5^11 because when multiplying numbers that have the same base but different exponents, you simply add the exponents. So, 5^7(5^4) = 5^7+4 or 5^11

OpenStudy (kendricklamar2014):

@Johnbc 5^2 and 5^8 is 5^10 not 5^11

OpenStudy (anonymous):

I am aware of that? I did not write that

OpenStudy (anonymous):

lol @KendrickLamar2014 that's a 3

OpenStudy (kendricklamar2014):

Oh lol

OpenStudy (anonymous):

But now that you understand the concept you can now explain that in your own words

OpenStudy (kendricklamar2014):

Can i do the third way 5^5 * 5^6?

OpenStudy (anonymous):

same concept just change the numbers

OpenStudy (kendricklamar2014):

but can I do that ^